• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Fatiha

    Log entry time 00:03:13 on August 20, 2006

    Entry number 180996

    keyword=Shift Summary

     4:00pm Started the shift with no beam. Called MCC for an update, they    
            are going to send cw to hall C, as soon as they are done with    
            it, they will give us good beam. 
     4:50pm Q1 tripped, tried to reset it from here... didn't work, asked for 
            controlled access. Paged Rob. 
     5:30pm Back from the access, MCC needs 10 minutes to cycle the Dipole. 
     5:45pm asekd for 20uA. 
     5:50pm TS11 crashed, reboot it and rebooted coda as well. 
     7:35pm Beam down, EP ion chamber trip.  MCC is working on this.
    10:20pm Q1 tripped again, asked for controlled access. 
    10:55pm Back from the access. 
    11:15pm Ion chamber problem again, MCC will need an access. 

    List of good runs:

    2935 LD2 1.9M, 20 uA 2936 LD2 1.87 M 20 uA, ended because beam went off. 2937 LD2 0.36 M 20 uA, ended because of a problem with the ion chamber.